OpenStudy (rawfeeds-):

What is the slope-intercept form of the function that contains the points (5, 6) and (9, 14)? y = x +

OpenStudy (rawfeeds-):

@Photon336

OpenStudy (anonymous):

ok so first, find the slope, do you know how to do that?

OpenStudy (rawfeeds-):

um yes one sec

OpenStudy (anonymous):

ok, so when you find the slope that will be your "m" in y=mx + b The y intercept (b) you will find with the slope

OpenStudy (rawfeeds-):

@DamienGannon slope=2

OpenStudy (anonymous):

you should get y=2x+b if you plug that in

OpenStudy (anonymous):

ok so now take 1 point (doesnt matter which) and plug the coordinates for x and y

OpenStudy (rawfeeds-):

ok can you do it

OpenStudy (anonymous):

ok so basically what you should get is 6=2(5) + b now what you have to do is multiple 2 by 5 and you should get 6=10 + b now you have to seperate b alone and you will get your answer

OpenStudy (anonymous):

in case you dont know how, it's decently simple, subtract 10 from both sides

OpenStudy (anonymous):

you should get -4 = b which is your y-intercept

OpenStudy (anonymous):

so plugging all the things you know in you get y=2x-4

OpenStudy (rawfeeds-):

so y=2x-4

OpenStudy (anonymous):

hope this helped

OpenStudy (anonymous):

yes